### v2

@@ 1 Introduction
Recent advances in generative modeling—particularly diffusion models [1, 2] and autoregressive approaches [3, 4]—have substantially enhanced the quality of synthesized content across image [5, 6], video [7, 8], and 3D contents [9, 10]. These breakthroughs have increasingly demonstrated the potential to extend beyond content generation into broader AI paradigms [11, 12, 13, 14, 15, 16, 17, 18, 19]. Despite this progress, most generative models are trained using surrogate objectives such as maximum likelihood estimation or reconstruction loss, which are often misaligned with human perception [20, 21], semantics [22], or physical plausibility [23]. As a result, generated outputs can exhibit artifacts such as motion inconsistency [24, 25], or structural errors [26].
Reinforcement learning (RL) [27] offers a principled framework for optimizing non-differentiable, human-aligned, or environment-driven objectives through interaction. Originally designed for control and decision-making, RL has emerged as a powerful complement to generative models, enabling long-horizon optimization, feedback-aware learning, and preference-based fine-tuning. In recent works [5, 28, 29], RL has been increasingly adopted to improve controllability, enhance realism, and align generation with complex and structured objectives across multiple modalities.
@@ 2 RL Revolution
Reinforcement learning (RL) [27] provides a general framework for sequential decision-making, where an agent learns to act by interacting with an environment to maximize long-term rewards. It is uniquely characterized by delayed feedback and the need to trade off exploration and exploitation. In the past decade, RL has demonstrated remarkable success in simulated domains such as video games [30, 31], robotics [32, 33, 34], and continuous control [35, 36], showcasing its potential under well-defined dynamics and reward structures [37, 30, 38]. Yet, real-world applications remain limited. Practical deployments often suffer from sample inefficiency [39, 40], reward specification challenges [41, 42], and poor generalization [43, 44]. The diversity of real-world environments [45, 43] further complicates the development of universally effective algorithms.
@@
Phase I: RL as solving Markov Decision Processes (MDPs) [46, 47] through trial-and-error [48, 27].
Phase II: A fragmentation into subfields (e.g., offline RL [49, 50], model-based RL [51, 52]) to overcome specific bottlenecks.
Phase III: A redefinition of RL’s core problem—shifting from policy optimization in known environments to learning environment dynamics [53] and aligning with human intent [42, 54].
Phase IV: The emergence of RL as a general substrate for decision-making [55], interfacing with planning [56], simulation [57], and generative modeling [28, 7].
@@ 2.1 RL as MDP Solvers
The first phase of RL was grounded in the formalism of Markov Decision Processes (MDPs), which define sequential decision-making under uncertainty through a set of states, actions, transition dynamics, and reward functions [46, 47, 58]. The goal was to compute an optimal policy that maximizes expected cumulative reward. Early RL efforts closely mirrored the dynamic programming solutions for MDPs, assuming either full or sampled access to the environment.
@@
Value-based methods were historically dominant. Temporal-Difference (TD) learning [59] laid the foundation by updating value estimates from sampled transitions, blending Monte Carlo evaluation with dynamic programming. Algorithms such as Q-learning [60] and SARSA [61] introduced off-policy and on-policy learning variants, while tabular implementations were standard for small-scale problems.
@@
Policy-based methods gained traction for their natural handling of stochastic policies and continuous action spaces. The REINFORCE algorithm [62] introduced Monte Carlo policy gradients using the likelihood ratio trick, albeit with high variance. This led to actor-critic architectures, which paired a learned value function (critic) to reduce the variance of policy updates.
@@
The integration of deep learning dramatically extended the reach of both paradigms. The Deep Q-Network (DQN) [30] approximated the Q-function with a convolutional neural network and achieved human-level performance on Atari games using raw pixel inputs. Key enhancements such as Double DQN [63], Dueling networks [64], and prioritized replay [65] were later unified in the Rainbow agent [66]. In parallel, scalable policy-based methods emerged. Asynchronous Advantage Actor-Critic (A3C)[67] introduced parallelism and advantage estimation; Trust Region Policy Optimization (TRPO)[68] and its practical variant Proximal Policy Optimization (PPO) [69] stabilized updates with theoretically grounded constraints, becoming standard in continuous control.
@@
These advances solidified deep RL as a powerful tool for solving complex MDPs, with notable successes in game playing [70], robotic manipulation [71], and simulated control [37]. However, the MDP-centric paradigm revealed several fundamental limitations that shaped the next phase of research. (1) Sample inefficiency. Most algorithms required millions of environment interactions to learn effective policies—tractable in simulation but prohibitive in domains like robotics or healthcare. This inefficiency arose from poor data reuse, slow policy updates, and especially inefficient exploration, which struggled in sparse-reward scenarios where agents failed to discover meaningful trajectories or assign credit over long horizons. (2) Limited generalization. RL agents were typically evaluated in the same environments they were trained on, leading to overfitting to visual or structural artifacts. Benchmarks like Procgen revealed that even small perturbations could significantly degrade performance, highlighting poor robustness and weak transferability. (3) Unrealistic assumptions. Classical RL methods often relied on fully observable, stationary environments with well-specified rewards and unrestricted interaction—conditions rarely met in practice. In many real-world tasks, interaction is costly or unsafe, data is offline or limited, and objectives are implicit or multi-faceted, making these assumptions untenable.
@@ 2.2 Emergence of Specialized Subfields
However, this paradigm introduces a central challenge: distributional shift between the behavior policy that collected the data and the target policy being learned. Without access to the environment, agents may assign erroneously high value to out-of-distribution actions, leading to extrapolation error and unstable training [49]. Recent methods address this through conservative value estimation or policy regularization. For example, CQL [49] penalizes Q-values for unseen actions to avoid overestimation, while BRAC [72] and TD3+BC [73] constrain the learned policy to remain close to the data distribution. IQL [74] avoids explicit constraints but modifies the Bellman update to implicitly favor in-distribution actions. These advances, along with growing theoretical understanding of bootstrapping error and pessimism, have made offline RL a promising solution for safe, sample-efficient learning from static logs—while also raising new challenges in data coverage, evaluation, and generalization.
@@
Multi-Agent RL: When multiple agents interact in the same environment, the single-agent MDP framework no longer holds because the environment becomes non-stationary from each agent’s perspective. This challenge led to the development of multi-agent RL (MARL) [75, 76, 77, 78], which addresses issues such as non-stationarity, cooperation, competition, and equilibrium strategies. Key contributions include MADDPG [79], which introduced centralized training with decentralized execution, allowing agents to condition their critics on global information during training while executing independently at test time. Other approaches, such as LOLA [80], incorporated opponent modeling by anticipating how an agent’s updates influence others’ learning. MARL expanded the classical RL framework to stochastic games and multi-agent MDPs, drawing on concepts from game theory to develop stable learning dynamics and policy solutions.
@@
Risk-Sensitive and Robust RL: Traditional RL optimizes the expected return, which may overlook performance variability or safety-critical failures. Risk-sensitive methods [81] address this by optimizing alternative criteria, such as minimizing variance or ensuring favorable worst-case outcomes. Distributional RL models [82, 83] the entire return distribution rather than its mean, enabling richer decision-making. Robust RL [84, 85, 86] further addresses uncertainty in dynamics or rewards, seeking policies that perform well under perturbations or adversarial changes. Techniques such as domain randomization [87] and adversarial training [86] have been used to enhance robustness and stress-test policies. These approaches reflect a shift from reward maximization toward more reliable and resilient objectives.
@@
Safe RL: Safe RL focuses on satisfying explicit safety constraints throughout training and deployment. It often formulates the problem as a constrained MDP, incorporating methods such as Lagrangian optimization [88], reward shaping [89], or safety shields [90, 91] to enforce limits on unsafe behavior. This is particularly critical in domains like autonomous driving [92] and robotics [93], where unsafe exploration can lead to catastrophic failures. Safe RL also intersects with exploration research, aiming to develop learning strategies that remain within safety boundaries while still enabling policy improvement.
@@ 2.3 Redefining the Problem — RL as Learning to Simulate, Optimize, and Align
During the late 2010s to mid-2020s, RL evolved from optimizing fixed reward functions in known environments to serving as a modular component within broader learning systems. This phase emphasized two key directions: first, using RL to optimize over models learned from data, such as simulators or reward functions [53, 52]; second, aligning agent behavior with nuanced human goals that cannot be easily captured by scalar rewards [42, 54]. The focus shifted from refining policy optimization algorithms to designing the outer structure of the learning task itself, including how preferences are modeled and how environment dynamics are inferred [94, 95]. RL became less about the optimizer and more about how the problem is formulated.
@@
RL from Human Feedback: RL from Human Feedback (RLHF) [42, 54] has become a high-impact application of RL, particularly in aligning large language models with human intentions. Rather than relying on hand-crafted reward functions, RLHF uses human preferences to define the objective. The typical pipeline involves collecting human demonstrations or rankings of model outputs [42, 94], training a reward model to predict these preferences, and then fine-tuning the model using RL to maximize this learned reward [54]. This transforms vague goals like “be helpful” into a concrete RL problem: the model generates responses, the reward is given by the preference model, and RL optimizes accordingly.
@@
The success of InstructGPT [54] demonstrated the effectiveness of this approach. A 1.3B parameter model fine-tuned via RLHF outperformed the original 175B GPT-3 in instruction following, while reducing harmful or irrelevant outputs. RLHF builds on earlier work such as preference-based RL [42], which showed that reward functions can be learned from human comparisons rather than manually specified. In high-dimensional spaces like language or vision, such feedback provides a scalable supervision signal. Importantly, the core challenge in RLHF lies not in the RL algorithm itself, but in designing robust reward models from subjective data [95]. This paradigm has since extended to domains like image generation and robotics, where human-aligned behavior is critical [94]. Overall, RLHF exemplifies a broader view of RL—as a flexible framework for aligning complex models with implicit goals through interaction and evaluative feedback.
@@
World Models and the Resurgence of Model-Based RL: Model-based RL has re-emerged as a promising solution to the sample inefficiency of traditional RL [57, 51]. By learning a model of environment dynamics from data, agents can simulate interactions internally, enabling policy training with significantly fewer real-world interactions. Early work such as World Models and PETS demonstrated that policies trained in learned simulators could transfer successfully to real environments [57, 51], highlighting the potential of model-based approaches.
@@
This direction was further advanced by Dreamer and its successors, which learned latent dynamics models and trained agents entirely within imagined trajectories [52, 96]. DreamerV2 [96] notably achieved competitive performance on the Atari benchmark using far fewer environment frames, demonstrating scalability to high-dimensional, vision-based tasks [96]. MuZero took a different approach by learning an implicit model to support planning via tree search, blending learning and planning without requiring a full transition model [53].
@@
Recent architectures such as DreamerV3 and Diffusion-QL integrate world models with generative and preference-aware objectives, illustrating a shift in focus from algorithmic innovation to task specification and system-level design [97, 98]. In this view, RL increasingly involves first learning the underlying Markov decision process—its dynamics and rewards—followed by planning or policy optimization within that learned model. This two-stage formulation offers greater flexibility and data efficiency, and has proven effective even when the learned model is imperfect. The result is a broader, more adaptable vision of RL, where modeling, simulation, and optimization are deeply intertwined.

### v3

@@ Abstract
00footnotetext: Survey homepage: https://visgenrlsurvey.liangyzh18.workers.dev/
@@ 1 Introduction
Recent advances in generative modeling—particularly diffusion models [yang2023diffusion, croitoru2023diffusion] and autoregressive approaches [xiong2024autoregressive, kaur2023autoregressive]—have substantially enhanced the quality of synthesized content across image [dhariwal2021diffusion, rombach2022high], video [rdpo, seedance], and 3D contents [ye2024dreamreward, lin2020modeling]. These breakthroughs have increasingly demonstrated the potential to extend beyond content generation into broader AI paradigms [liang2019vrr, AIFlowReport, chen2024hydra, PN, VPN, PiNDA, PiNI, PiNGDA, MiN]. Despite this progress, most generative models are trained using surrogate objectives such as maximum likelihood estimation or reconstruction loss, which are often misaligned with human perception [zhang2018unreasonable, esser2021taming], semantics [radford2021learning], or physical plausibility [ho2022video]. As a result, generated outputs can exhibit artifacts such as motion inconsistency [villegas2017learning, yan2021videogpt], or structural errors [mescheder2018training].
Reinforcement learning (RL) [sutton1998reinforcement] offers a principled framework for optimizing non-differentiable, human-aligned, or environment-driven objectives through interaction. Originally designed for control and decision-making, RL has emerged as a powerful complement to generative models, enabling long-horizon optimization, feedback-aware learning, and preference-based fine-tuning. In recent works [dhariwal2021diffusion, dancegrpo, zhou2024flowgrpo], RL has been increasingly adopted to improve controllability, enhance realism, and align generation with complex and structured objectives across multiple modalities.
@@ 2 RL Revolution
Reinforcement learning (RL) [sutton1998reinforcement] provides a general framework for sequential decision-making, where an agent learns to act by interacting with an environment to maximize long-term rewards. It is uniquely characterized by delayed feedback and the need to trade off exploration and exploitation. In the past decade, RL has demonstrated remarkable success in simulated domains such as video games [mnih2015human, shao2019survey], robotics [ibarz2021train, zhao2020sim, kalashnikov2018scalable], and continuous control [dulac2021challenges, lillicrap2015continuous], showcasing its potential under well-defined dynamics and reward structures [tassa2018deepmind, mnih2015human, wang2024distributionally]. Yet, real-world applications remain limited. Practical deployments often suffer from sample inefficiency [mohanty2021measuring, kaiser2019model], reward specification challenges [ng1999policy, christiano2017deep], and poor generalization [zhang2018study, arjovsky2020out]. The diversity of real-world environments [dulac2019challenges, zhang2018study] further complicates the development of universally effective algorithms.
@@
Phase I: RL as solving Markov Decision Processes (MDPs) [bellman2015applied, puterman2014markov] through trial-and-error [sutton1991dyna, sutton1998reinforcement].
Phase II: A fragmentation into subfields (e.g., offline RL [kumar2020conservative, fujimoto2019off], model-based RL [chua2018deep, hafner2019dream]) to overcome specific bottlenecks.
Phase III: A redefinition of RL’s core problem—shifting from policy optimization in known environments to learning environment dynamics [schrittwieser2020mastering] and aligning with human intent [christiano2017deep, ouyang2022training].
Phase IV: The emergence of RL as a general substrate for decision-making [hafner2023mastering], interfacing with planning [silver2018general], simulation [ha2018worldmodels], and generative modeling [dancegrpo, rdpo].
@@ 2.1 RL as MDP Solvers
The first phase of RL was grounded in the formalism of Markov Decision Processes (MDPs), which define sequential decision-making under uncertainty through a set of states, actions, transition dynamics, and reward functions [bellman2015applied, puterman2014markov, bellman1957dp]. The goal was to compute an optimal policy that maximizes expected cumulative reward. Early RL efforts closely mirrored the dynamic programming solutions for MDPs, assuming either full or sampled access to the environment.
@@
Value-based methods were historically dominant. Temporal-Difference (TD) learning [sutton1988td] laid the foundation by updating value estimates from sampled transitions, blending Monte Carlo evaluation with dynamic programming. Algorithms such as Q-learning [watkins1992q] and SARSA [rummery1994line] introduced off-policy and on-policy learning variants, while tabular implementations were standard for small-scale problems.
@@
Policy-based methods gained traction for their natural handling of stochastic policies and continuous action spaces. The REINFORCE algorithm [williams1992simple] introduced Monte Carlo policy gradients using the likelihood ratio trick, albeit with high variance. This led to actor-critic architectures, which paired a learned value function (critic) to reduce the variance of policy updates.
@@
The integration of deep learning dramatically extended the reach of both paradigms. The Deep Q-Network (DQN) [mnih2015human] approximated the Q-function with a convolutional neural network and achieved human-level performance on Atari games using raw pixel inputs. Key enhancements such as Double DQN [hasselt2016deep], Dueling networks [wang2016dueling], and prioritized replay [schaul2016prioritized] were later unified in the Rainbow agent [hessel2018rainbow]. In parallel, scalable policy-based methods emerged. Asynchronous Advantage Actor-Critic (A3C)[mnih2016asynchronous] introduced parallelism and advantage estimation; Trust Region Policy Optimization (TRPO)[schulman2015trust] and its practical variant Proximal Policy Optimization (PPO) [schulman2017proximal] stabilized updates with theoretically grounded constraints, becoming standard in continuous control.
@@
These advances solidified deep RL as a powerful tool for solving complex MDPs, with notable successes in game playing [silver2016mastering], robotic manipulation [levine2016end], and simulated control [tassa2018deepmind]. However, the MDP-centric paradigm revealed several fundamental limitations that shaped the next phase of research. (1) Sample inefficiency. Most algorithms required millions of environment interactions to learn effective policies—tractable in simulation but prohibitive in domains like robotics or healthcare. This inefficiency arose from poor data reuse, slow policy updates, and especially inefficient exploration, which struggled in sparse-reward scenarios where agents failed to discover meaningful trajectories or assign credit over long horizons. (2) Limited generalization. RL agents were typically evaluated in the same environments they were trained on, leading to overfitting to visual or structural artifacts. Benchmarks like Procgen revealed that even small perturbations could significantly degrade performance, highlighting poor robustness and weak transferability. (3) Unrealistic assumptions. Classical RL methods often relied on fully observable, stationary environments with well-specified rewards and unrestricted interaction—conditions rarely met in practice. In many real-world tasks, interaction is costly or unsafe, data is offline or limited, and objectives are implicit or multi-faceted, making these assumptions untenable.
@@ 2.2 Emergence of Specialized Subfields
However, this paradigm introduces a central challenge: distributional shift between the behavior policy that collected the data and the target policy being learned. Without access to the environment, agents may assign erroneously high value to out-of-distribution actions, leading to extrapolation error and unstable training [kumar2020conservative]. Recent methods address this through conservative value estimation or policy regularization. For example, CQL [kumar2020conservative] penalizes Q-values for unseen actions to avoid overestimation, while BRAC [wu2019behavior] and TD3+BC [fujimoto2021minimalist] constrain the learned policy to remain close to the data distribution. IQL [kostrikov2021offline] avoids explicit constraints but modifies the Bellman update to implicitly favor in-distribution actions. These advances, along with growing theoretical understanding of bootstrapping error and pessimism, have made offline RL a promising solution for safe, sample-efficient learning from static logs—while also raising new challenges in data coverage, evaluation, and generalization.
@@
Multi-Agent RL: When multiple agents interact in the same environment, the single-agent MDP framework no longer holds because the environment becomes non-stationary from each agent’s perspective. This challenge led to the development of multi-agent RL (MARL) [canese2021multi, zhang2021multi, bucsoniu2010multi, foerster2018counterfactual], which addresses issues such as non-stationarity, cooperation, competition, and equilibrium strategies. Key contributions include MADDPG [lowe2017multi], which introduced centralized training with decentralized execution, allowing agents to condition their critics on global information during training while executing independently at test time. Other approaches, such as LOLA [foerster2017learning], incorporated opponent modeling by anticipating how an agent’s updates influence others’ learning. MARL expanded the classical RL framework to stochastic games and multi-agent MDPs, drawing on concepts from game theory to develop stable learning dynamics and policy solutions.
@@
Risk-Sensitive and Robust RL: Traditional RL optimizes the expected return, which may overlook performance variability or safety-critical failures. Risk-sensitive methods [chow2015risk] address this by optimizing alternative criteria, such as minimizing variance or ensuring favorable worst-case outcomes. Distributional RL models [bellemare2017distributional, dabney2018distributional] the entire return distribution rather than its mean, enabling richer decision-making. Robust RL [pinto2017robust, vinitsky2020robust, kamalaruban2020robust] further addresses uncertainty in dynamics or rewards, seeking policies that perform well under perturbations or adversarial changes. Techniques such as domain randomization [tobin2017domain] and adversarial training [kamalaruban2020robust] have been used to enhance robustness and stress-test policies. These approaches reflect a shift from reward maximization toward more reliable and resilient objectives.
@@
Safe RL: Safe RL focuses on satisfying explicit safety constraints throughout training and deployment. It often formulates the problem as a constrained MDP, incorporating methods such as Lagrangian optimization [achiam2017constrained], reward shaping [garcia2015comprehensive], or safety shields [alshiekh2018safe, elsayed2021safe] to enforce limits on unsafe behavior. This is particularly critical in domains like autonomous driving [zhao2024survey] and robotics [berkenkamp2017safe], where unsafe exploration can lead to catastrophic failures. Safe RL also intersects with exploration research, aiming to develop learning strategies that remain within safety boundaries while still enabling policy improvement.
@@ 2.3 Redefining the Problem — RL as Learning to Simulate, Optimize, and Align
During the late 2010s to mid-2020s, RL evolved from optimizing fixed reward functions in known environments to serving as a modular component within broader learning systems. This phase emphasized two key directions: first, using RL to optimize over models learned from data, such as simulators or reward functions [schrittwieser2020mastering, hafner2019dream]; second, aligning agent behavior with nuanced human goals that cannot be easily captured by scalar rewards [christiano2017deep, ouyang2022training]. The focus shifted from refining policy optimization algorithms to designing the outer structure of the learning task itself, including how preferences are modeled and how environment dynamics are inferred [ziegler2019fine, bai2022training]. RL became less about the optimizer and more about how the problem is formulated.
@@
RL from Human Feedback: RL from Human Feedback (RLHF) [christiano2017deep, ouyang2022training] has become a high-impact application of RL, particularly in aligning large language models with human intentions. Rather than relying on hand-crafted reward functions, RLHF uses human preferences to define the objective. The typical pipeline involves collecting human demonstrations or rankings of model outputs [christiano2017deep, ziegler2019fine], training a reward model to predict these preferences, and then fine-tuning the model using RL to maximize this learned reward [ouyang2022training]. This transforms vague goals like “be helpful” into a concrete RL problem: the model generates responses, the reward is given by the preference model, and RL optimizes accordingly.
@@
The success of InstructGPT [ouyang2022training] demonstrated the effectiveness of this approach. A 1.3B parameter model fine-tuned via RLHF outperformed the original 175B GPT-3 in instruction following, while reducing harmful or irrelevant outputs. RLHF builds on earlier work such as preference-based RL [christiano2017deep], which showed that reward functions can be learned from human comparisons rather than manually specified. In high-dimensional spaces like language or vision, such feedback provides a scalable supervision signal. Importantly, the core challenge in RLHF lies not in the RL algorithm itself, but in designing robust reward models from subjective data [bai2022training]. This paradigm has since extended to domains like image generation and robotics, where human-aligned behavior is critical [ziegler2019fine]. Overall, RLHF exemplifies a broader view of RL—as a flexible framework for aligning complex models with implicit goals through interaction and evaluative feedback.
@@
World Models and the Resurgence of Model-Based RL: Model-based RL has re-emerged as a promising solution to the sample inefficiency of traditional RL [ha2018worldmodels, chua2018deep]. By learning a model of environment dynamics from data, agents can simulate interactions internally, enabling policy training with significantly fewer real-world interactions. Early work such as World Models and PETS demonstrated that policies trained in learned simulators could transfer successfully to real environments [ha2018worldmodels, chua2018deep], highlighting the potential of model-based approaches.
@@
This direction was further advanced by Dreamer and its successors, which learned latent dynamics models and trained agents entirely within imagined trajectories [hafner2019dream, hafner2020dreamerv2]. DreamerV2 [hafner2020dreamerv2] notably achieved competitive performance on the Atari benchmark using far fewer environment frames, demonstrating scalability to high-dimensional, vision-based tasks [hafner2020dreamerv2]. MuZero took a different approach by learning an implicit model to support planning via tree search, blending learning and planning without requiring a full transition model [schrittwieser2020mastering].
@@
Recent architectures such as DreamerV3 and Diffusion-QL integrate world models with generative and preference-aware objectives, illustrating a shift in focus from algorithmic innovation to task specification and system-level design [hafner2023dreamerv3, wang2022diffusion]. In this view, RL increasingly involves first learning the underlying Markov decision process—its dynamics and rewards—followed by planning or policy optimization within that learned model. This two-stage formulation offers greater flexibility and data efficiency, and has proven effective even when the learned model is imperfect. The result is a broader, more adaptable vision of RL, where modeling, simulation, and optimization are deeply intertwined.